OpenStudy (anonymous):

Help: A certain recipe asks for 2/9 cup of granola. How much is needed to make 4/3 of a recipe?

OpenStudy (anonymous):

(2/9)x=4/3 ----> x=(4/3)(9/2)=6 cups

OpenStudy (anonymous):

thankyou :)